OpenStudy (vortish):

two angles are supplementary one is 30 little circle above the 30 more than four times the other find the measure of the angles

OpenStudy (vortish):

what is the measure of the smaller angle and what is measure of the larger angle

OpenStudy (mathstudent55):

Little circle above is the symbol for degree or degrees which is a unit used to measure angles. So you're trying to say: "Two angles are supplementary. One is 30 degrees more than four times the other, Find the measure of the angles." Is this correct?

OpenStudy (vortish):

yes

OpenStudy (mathstudent55):

Then let x equal measure of smaller angle. The other angle is 30 degrees more than 4 times x. How would you represent that with x?

OpenStudy (vortish):

30 + 4x

OpenStudy (mathstudent55):

Correct. Do you know what supplementary angles means?

OpenStudy (vortish):

no

OpenStudy (mathstudent55):

Two angles are supplementary if the sum of their measures is 180 degrees.

OpenStudy (mathstudent55):

Now you know three things: The measure of smaller angle is x The measure of larger angle is 30 + 4x The sum of the measures is 180 Can you represent that sum using the info above?

OpenStudy (vortish):

30+4x=180

OpenStudy (vortish):

if that is true then 4x=150

OpenStudy (mathstudent55):

almost, one angle is x the other 30 + 4x the sum of those two things equals 180

OpenStudy (mathstudent55):

30 + 4x is the larger angle. That does not equal 180

OpenStudy (vortish):

so it woud be angle one = x angle two = 30+4x

OpenStudy (vortish):

so x +30 +4x = 180

OpenStudy (mathstudent55):

correct

OpenStudy (vortish):

30 +5x=180 5x=150 x = 30

OpenStudy (mathstudent55):

Correct. That's the measure of the smaller angle. Now what's the measure of the larger one?

OpenStudy (vortish):

so one angle 150 and the other is 30

OpenStudy (mathstudent55):

you got it

OpenStudy (vortish):

thanks math you made me do the work but gave me the steps to do it i work and retain better that way apperciate it

OpenStudy (mathstudent55):

you're welcome